# thanks: Authors of equal contributionsthanks: Authors of equal contributions

### v2

@@ Introduction.
We utilize the LCU algorithm to propagate the bosonic and fermionic DQME-SQ exemplified with spin-boson model and single-impurity Anderson model, respectively. To exhibit the feasibility of our framework, we employ the Aer simulator of Qiskit Qiskit to propagate the DQME-SQ. Within each time step, we directly extract the target state, bypassing the need for extensive probabilistic measurements required on actual quantum computing platforms.
@@
As shown in Fig. 3, the simulated time evolution results obtained from DQME-SQ and HEOM exhibit perfect agreement, where the HEOM method utilizes classical computation algorithms Li23SM . Notably, Fig. 3 demonstrates that non-Markovian effects are more pronounced at lower temperatures in both bosonic and fermionic environments. With all other parameters unchanged, a lower temperature leads to more pronounced oscillations in the short-time regime and slower relaxation in the long-time regime, due to longer environmental memory.
Due to limited computational resources at our disposal, the digital quantum simulations presented above are constrained to minimal models. However, the utilized quantum algorithm will become universally feasible with significant advancements in future quantum computing technologies.
Summary and perspective.
Towards the quantum simulation of non-Markovian open quantum dynamics, this paper presents the DQME-SQ theory, a universal and exact theory that enables quantum simulation of dissipative dynamics in both bosonic and fermionic environments. By translating the non-Markovian memory content into a compact description of statistical quasi-particles (dissipatons), the DQME-SQ theory provides a physical interpretation of OQS that is particularly suited for quantum computation, especially in fermionic environments. This framework provides extensive universality and flexibility for advancing quantum simulation in OQS, enhancing studies of nonequilibrium dynamics and thermodynamics by elucidating the roles of non-Markovian influences in complex OQS Cle101155 ; Soa14825 ; Coh15266802 ; Har201184 .

### v3

@@ V Numerical demonstration through digital quantum simulation
To demonstrate the feasibility of the DQME-SQ approach, we employ the LCU-Trotter algorithm to propagate the quantum dissipative dynamics for two exemplary models: the spin-boson model for the bosonic case, and the single-impurity Anderson model for the fermionic case. For this purpose, we utilize the Aer simulator of Qiskit Qiskit to propagate the DQME-SQ. Within each time step, we directly extract the target state, thus bypassing the need for extensive probabilistic measurements that would be required on actual quantum computing platforms.
@@
As shown in Fig. 3, the simulated time evolution results obtained from DQME-SQ and HEOM exhibit perfect agreement, where the HEOM method utilizes classical computation algorithms. Notably, Fig. 3 demonstrates that non-Markovian effects are more pronounced at lower temperatures in both bosonic and fermionic environments. With all other parameters unchanged, a lower temperature leads to more pronounced oscillations in the short-time regime and slower relaxation in the long-time regime, due to longer environmental memory.
Simulations of more complex OQSs are demonstrated in the Appendix C. These additional examples further exhibit the exactness and universality of the DQME-SQ theory. Due to limited computational resources at our disposal, the digital quantum simulations presented above are constrained to minimal models. When implementing the DQME-SQ approach on realistic NISQ devices, the errors and measurement costs arising from the unitary decomposition approximations will need to be carefully considered and addressed. Techniques such as those proposed in Refs. Sch21270503 ; Sch22023216 can be employed to mitigate these issues. However, as quantum computing technologies continue to advance significantly in the future, the utilized quantum algorithm, including the LCU-Trotter method, will become universally feasible for simulating more complex OQSs.
VI Summary and perspective
Towards the quantum simulation of non-Markovian open quantum dynamics, this paper presents the DQME-SQ theory, a universal and exact theory that enables quantum simulation of dissipative dynamics in both bosonic and fermionic environments. It captures the full non-Markovian nature of open quantum systems, provided that the environments follow Gaussian statistics. By translating the non-Markovian memory content into a compact description of statistical quasi-particles (dissipatons), the DQME-SQ theory provides a physical interpretation of OQSs that is particularly suited for quantum computation, especially in fermionic environments. This framework provides extensive universality and flexibility for advancing quantum simulation in OQSs, enhancing studies of nonequilibrium dynamics and thermodynamics by elucidating the roles of non-Markovian influences in complex OQSs Cle101155 ; Soa14825 ; Coh15266802 ; Har201184 . Moreover, it is also intriguing to explore the connection between the DQME-SQ and other recently developed methods, such as those based on the ensemble of trajectories Hea19022109 and the density matrix purification Sch22_arXiv_2207_07112 . These methods may share some useful techniques that could be beneficial for the quantum simulation of open quantum dynamics.